# Galois Correspondence for Partial Groupoid Actions

### v1

@@ 4.1. Orthogonalization
It only takes us to observe that we can realize $E=\prod_{e\in\mathcal{G}_{0}}A_{e}$ as an ideal of $B$ and that $\varepsilon$ coincides with the standard restriction of $\beta$ to $E$ as defined in [2, Section 1], that is, $E_{e}=B_{e}\cap E$, for all $e\in\mathcal{G}_{0}$, $E_{g}=E_{r(g)}\cap\beta_{g}(E_{d(g)})$, for all $g\in\mathcal{G}$, and $\varepsilon_{g}=\beta_{g}|_{E_{g^{-1}}}$. ∎
@@
Let $\mathcal{G}=\{g,g^{-1},r(g),d(g)\}$, $A=Re_{1}\oplus Re_{2}\oplus Re_{3}$, where $R\neq 0$ is a commutative ring and the $e_{i}$’s are pairwise orthogonal, central idempotents such that $e_{1}+e_{2}+e_{3}=1_{A}$.
@@ 4.2. Strongly $\alpha$-Galois Extension
Let $a\in A^{\alpha|_{\mathcal{H}}}$. Given $h\in\mathcal{H}$,
@@
For the reverse one, consider $a\in A$ such that $\varphi(a)\in E^{\varepsilon|_{\mathcal{H}}}$. Then, given $h\in\mathcal{H}$, we have that $\varepsilon_{h}(\varphi(a)\varphi_{d(h)}(1_{h^{-1}}))=\varphi(a)\varphi_{r(h)}(1_{h})$. However
@@
Therefore $\varphi_{r(h)}(\alpha_{h}(a1_{h^{-1}}))=\varphi_{r(h)}(a1_{h})$. Since $\varphi_{r(h)}$ is a monomorphism, it follows that $\alpha_{h}(a1_{h^{-1}})=a1_{h}$, for all $h\in\mathcal{H}$, that is, $a\in A^{\alpha|_{\mathcal{H}}}$. ∎
@@
Let $\alpha$ be a partial action of the groupoid $\mathcal{G}$ on a ring $A$. Consider $\varepsilon$ the orthogonalization of $\alpha$ on a ring $E$. We say that $\alpha$ is a strongly Galois partial action if:
@@
If $C\in\mathsf{A}$, then $C=\varphi^{-1}(E^{\varepsilon|_{\mathcal{H}}}\cap\varphi(A))$, for some $\mathcal{H}\in\mathsf{W}$. By Proposition 4.8, $C=\varphi^{-1}(\varphi(A^{\alpha|_{\mathcal{H}}}))=A^{\alpha|_{\mathcal{H}}}$. Now, consider $\{x_{i},y_{i}\}_{i=1}^{n}$ a partial $\alpha$-Galois coordinate system for $A$ over $A^{\alpha}$, that is,
@@
Now we can prove the main result of this subsection.
@@
    Theorem 4.11 gives us a generalization of the Galois correspondences that already exist in the literature, but it is still restrictive, once it asks that the partial action is strongly Galois, and not all partial actions over Galois extensions belongs to this class, as we can see in the next example.
@@
Let $\mathcal{G}=\{g,f_{1}\}\cup\{h,f_{2}\}$ the groupoid given by the disjoint union of two groups isomorphic to $\mathbb{Z}_{2}$, that is, $g^{2}=f_{1}=f_{1}^{2}$ and $h^{2}=f_{2}=f_{2}^{2}$. Consider a commutative ring $R$ and $A=Re_{1}\oplus Re_{2}$, with $e_{1},e_{2}$ orthogonal idempotents whose sum is $1_{A}$. Define $A_{k}=A$, for all $k\in\mathcal{G}$,
@@ 4.3. Weakening the hypothesis
It is true that $A^{\beta|_{\mathcal{H}}}\subseteq A^{\beta|_{\mathcal{H}_{1}}}$, since $\mathcal{H}_{1}\subseteq\mathcal{H}$. For the reverse inclusion, let $a\in A^{\beta|_{\mathcal{H}_{1}}}=A^{\beta|_{\mathcal{H}_{2}}}$. Let $h\in\mathcal{H}$. We have that $h=h_{1}\cdots h_{n}$, for some positive integer $n$ and $h_{1},\ldots,h_{n}\in\mathcal{H}_{1}\cup\mathcal{H}_{2}$ with $d(h_{i})=r(h_{i+1})$, for all $11\leq i\leq n-$ 1. Then
@@
Let $\mathcal{G}$ be a finite groupoid acting on a commutative ring $A$ via a unital global action $\beta$. Consider $\varepsilon$ the orthogonalization of $\beta$ on a ring $E$. Assume that $A_{g}\neq 0$, for all $g\in\mathcal{G}_{0}$, and that $A$ is a $\beta$-Galois extension of $A^{\beta}$. Then there is a one-to-one correspondence between the sets $\mathsf{W}^{max}$ and $\mathsf{A}$ given by $\mathsf{W}^{max}\ni\mathcal{H}\mapsto A^{\beta|_{\mathcal{H}}}$ with inverse $\mathsf{A}\ni C\mapsto(\mathcal{G}_{\varphi(C)})^{max}$.
@@
So, by Claim 1 and Claim 2, there is a one-to-one correspondence between the sets $\mathsf{W}^{max}$ and $\mathsf{A}$.

### v2

@@ 4.1. Orthogonalization
It only takes to observe that we can realize $E=\prod_{e\in\mathcal{G}_{0}}A_{e}$ as an ideal of $B$ and that $\varepsilon$ coincides with the standard restriction of $\beta$ to $E$ as defined in [2, Section 1], that is, $E_{e}=B_{e}\cap E$, for all $e\in\mathcal{G}_{0}$, $E_{g}=E_{r(g)}\cap\beta_{g}(E_{d(g)})$, for all $g\in\mathcal{G}$, and $\varepsilon_{g}=\beta_{g}|_{E_{g^{-1}}}$. ∎
@@
Consider $\mathcal{G}=\{g,g^{-1},r(g),d(g)\}$ and $A=Re_{1}\oplus Re_{2}\oplus Re_{3}$, where $R\neq 0$ is a commutative ring and the $e_{i}$’s are pairwise orthogonal, central idempotents such that $e_{1}+e_{2}+e_{3}=1_{A}$.
@@ 4.2. Strongly $\alpha$-Galois Extension
Take $a\in A^{\alpha|_{\mathcal{H}}}$. Given $h\in\mathcal{H}$,
@@
For the reverse, consider $a\in A$ such that $\varphi(a)\in E^{\varepsilon|_{\mathcal{H}}}$. Then, given $h\in\mathcal{H}$, we have that $\varepsilon_{h}(\varphi(a)\varphi_{d(h)}(1_{h^{-1}}))=\varphi(a)\varphi_{r(h)}(1_{h})$. However
@@
Therefore, $\varphi_{r(h)}(\alpha_{h}(a1_{h^{-1}}))=\varphi_{r(h)}(a1_{h})$. Since $\varphi_{r(h)}$ is a monomorphism, it follows that $\alpha_{h}(a1_{h^{-1}})=a1_{h}$, for all $h\in\mathcal{H}$, that is, $a\in A^{\alpha|_{\mathcal{H}}}$. ∎
@@
Let $\alpha$ be a partial action of the groupoid $\mathcal{G}$ on a ring $A$. Consider the orthogonalization $\varepsilon$ of $\alpha$ on a ring $E$. We say that $\alpha$ is a strongly Galois partial action if:
@@
If $C\in\mathsf{A}$, then $C=\varphi^{-1}(E^{\varepsilon|_{\mathcal{H}}}\cap\varphi(A))$, for some $\mathcal{H}\in\mathsf{W}$. By Proposition 4.8, $C=\varphi^{-1}(\varphi(A^{\alpha|_{\mathcal{H}}}))=A^{\alpha|_{\mathcal{H}}}$. Now, consider a partial $\alpha$-Galois coordinate system $\{x_{i},y_{i}\}_{i=1}^{n}$ for $A$ over $A^{\alpha}$, that is,
@@
Now we can prove the main result of this subsection. From now on, assume that the ring $A$ is commutative.
@@
    Theorem 4.11 gives a generalization of the Galois correspondences that already exist in the literature, but it is still restrictive, once it requires that the partial action is strongly Galois, and not all partial actions over Galois extensions belongs to this class, as we can see in the next example.
@@
Let $\mathcal{G}=\{g,f_{1}\}\cup\{h,f_{2}\}$ be the groupoid given by the disjoint union of two groups isomorphic to $\mathbb{Z}_{2}$, that is, $g^{2}=f_{1}=f_{1}^{2}$ and $h^{2}=f_{2}=f_{2}^{2}$. Consider a commutative ring $R$ and $A=Re_{1}\oplus Re_{2}$, with $e_{1},e_{2}$ orthogonal idempotents whose sum is $1_{A}$. Define $A_{k}=A$, for all $k\in\mathcal{G}$,
@@ 4.3. Weakening the hypothesis
It is true that $A^{\beta|_{\mathcal{H}}}\subseteq A^{\beta|_{\mathcal{H}_{1}}}$, since $\mathcal{H}_{1}\subseteq\mathcal{H}$. For the reverse inclusion, consider $a\in A^{\beta|_{\mathcal{H}_{1}}}=A^{\beta|_{\mathcal{H}_{2}}}$. Take $h\in\mathcal{H}$. We have that $h=h_{1}\cdots h_{n}$, for some positive integer $n$ and $h_{1},\ldots,h_{n}\in\mathcal{H}_{1}\cup\mathcal{H}_{2}$ with $d(h_{i})=r(h_{i+1})$, for all $11\leq i\leq n-$ 1. Then
@@
Let $\mathcal{G}$ be a finite groupoid acting on a commutative ring $A$ via a unital global action $\beta$. Consider the orthogonalization $\varepsilon$ of $\beta$ on a ring $E$. Assume that $A_{g}\neq 0$, for all $g\in\mathcal{G}_{0}$, and that $A$ is a $\beta$-Galois extension of $A^{\beta}$. Then there is a one-to-one correspondence between the sets $\mathsf{W}^{max}$ and $\mathsf{A}$ given by $\mathsf{W}^{max}\ni\mathcal{H}\mapsto A^{\beta|_{\mathcal{H}}}$ with inverse $\mathsf{A}\ni C\mapsto(\mathcal{G}_{\varphi(C)})^{max}$.
@@
So, by Claim 1 and Claim 2, there is a one-to-one correspondence between the sets $\mathsf{W}^{max}$ and $\mathsf{A}$. ∎